OpenStudy (anonymous):

Allan and Dave bowl together and their combined total score for one game was 375 points. Allan’s score was 60 less than twice Dave’s. What were their scores? Which is a system of equations to model the problem if x represents Dave’s score and y represents Allan’s score?

OpenStudy (anonymous):

answers could be x + y = 60 y = 2x – 375 x + y = 375 y = 2x – 60 x + y = 375 y = 2x + 60 x – y = 375 y = 2x – 60

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

y=2x-60 x+y=375

OpenStudy (anonymous):

thank you so much! @quixoticideals

OpenStudy (anonymous):

Sure. :)